A 80 turns closed circular coil having a radius of 20 cm is placed perpendicular to uniform magnetic field of intensity 50 Oe. Determine the amount of charge flowing through the coil, if the coil is rotated through 180° with respect to any diameter.

Resistance of the coil is given 2Ω.


1
Expert's answer
2021-02-18T18:39:21-0500

"H=\frac{B}{\mu_0},"

"B=\frac{\mu_0 n I}{2\pi r},"

"H=\frac{nI}{2\pi r},"

"I=\frac{2\pi rH}{n},"

"I=\frac{2\cdot 3.14\cdot 0.2\cdot 50}{80}=0.785 ~A
."